a bus travels for 45 minutes at 40 mph. how far does he go?

Answer by partha_ban(41)

45 minutes = 45 / 60 hours = 3/4 hours
Speed of the bus = 40 mph meaning in 1 hour the bus travells = 40 miles
Therefore, distance travelled in 3/4 hours = 40*3/4 miles = 30 miles